Q5 3 Points What is the maximum value of 22/(23 – 15) for |z< 2? O 2/15 0 1/4 04/7 O 47
Answer #1

Solution 3 The maximum value of اریک for 12152 = 2 let f(z) = 2 we see for maxima 22 - (کاتی 22 ازيگ و 30 یا درfor f1 (2) = 0 - - 2 C30 +23) = 0 = 0, -(30) critical points 2 Since maxima for 12152 that is 2. Z = 2 -(30)3 2-2 ffczi) ŽE